RC Meeting, Wednesday, September 15, 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings at 10:00, remote

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Tuesday Status

  • No beam until roughly mid swing shift due to RF separator work
    • This will continue (and hopefully finish) today
  • Lots of other work done though
  • Brad cycled the HGC gas see the log entry, thanks Brad!
  • Ali updated the report file output, things look more realistic now as noted in his log entry
    • He also improved the efficiency of the physics replays, should be ~10% faster now - log entry
    • Preliminary detector efficiencies also added - log entry
  • Carlos checked on the BCM calibrations and implemented them - log entry
  • I ran an updated SHMS Calorimeter calibration log entry
    • This doesn't seem to have percolated through to some replays and/or it is a bit off
    • Jacob and Simona are investigating
  • I re-analysed an earlier setting and found that our improvements to various param files have helped a lot - log entry
  • Simona was taking a closer look at the NGC following some quick analysis/checks I'd done in the morning, something look a bit odd log entry
    • It looks like one side of the NGC sees drastically less light than the other
  • Nearing the end of this kinematic point
    • Nathan and the shift crew overnight rotated out to our central (SHMS 8.26) setting - log entry
    • Not much longer at this angle once we get beam back, then we rotate out another 2 degrees to our left setting
    • Running at 15uA on LD2, rates look ok as posted in this log entry
    • See the status page for more
  • Thanks to everyone who's helped out on analysis this week, Simona, Carlos, Jacob, Richard, Ali, Nathan, Junaid, we're in much better shape this week than we were last week!

Current Status

  • Down due to Hall D startup work again
  • As far as I know, should expect beam in swing shift again

Current Run Plan

  • Carry on with current setting once we get beam back
  • Finish off the centre and left LD2- Q2=1.60, W=3.08 settings
  • HeeP Singles next
    • This will be a fairly intense block of shifts!

Current Issues

  • The HGC now looks great as mentioned... but now the NGC looks strange
    • Seeing next to nothing on one side of the fp
    • See this log entry

Planned Accesses to the Hall

  • We have no beam till ~ swing. I did not schedule a sweep, but if you want to go in for some reason, an escorted access should be possible